OpenStudy (anonymous):

\[\frac{ \sqrt{12} }{ 2 }\] give answer in form \[ a\sqrt{b}\]

OpenStudy (anonymous):

\[\sqrt{12}/2\] = \[\sqrt{2 \times 2 \times 3}/2\] = \[2\sqrt{3}/2\] = \[1\sqrt{3}\]